Find the following probabilities. Use drawings to assist when necessary. Round all probabilities to 4 decimal places.

P(z > 1.94)

P(z < -2.16)

P(-0.75 < z < 1.86)
Answer #1

Solution :

Using standard normal table,

P(z > 1.94) = 1 - P(z < 1.94) = 1 - 0.9738 = 0.0262

P(z < -2.16) = 0.0154

P(-0.75 < z < 1.86) = P(z < 1.86) - P(z < -0.75) = 0.9686 - 0.2266 = 0.7420